what mass of water can be heated from 25°c to 50°c by the addition of 2825 j of heat


1
Expert's answer
2021-10-26T05:34:03-0400

"T_1 = 25 \;°C \\

T_2 = 50 \;°C \\

Q = 2825 \;J \\

Q = mcΔT \\

c=4.186 \;J/g°C \\

m = \frac{Q}{cΔT} \\

m = \frac{2825}{4.186(50-25)} = 26.99 \;g"

Answer: 26.99 g